What is the pressure (in atm and mmHg) of 5.00 g of oxygen gas at 35oC that is

contained in a 255 mL container?

Can u please show dymentional analysis?
Answer #1

We can use the ideal gas equation,

P V = n R T

Where,

P = Pressure , V = Volume , n = No. of moles, R =- gas constant , T = temp in kelvin

We require number of moles of oxygen = Weight / mol wt

= 5 / 16 = 0.31

P x 255 = 0.31 x 8.314 x 308 [ 35 o C + 273 = 308 K ]

Therefore

P = 793.82 / 255

= 3.11 atm

= 2363.59 mmHg [ Since 1 atm = 760 mmHg, therefore 3.11 atm = 3.11 x 760 = 2363.59 mmHg ]